Ellen wishes to mix candy worth $1.15 per pound with candy worth $2.87 per pound to form 20 pounds of a mixture worth $1.58 per pound. How many pounds of the more expensive candy should she use?
Answer by stanbon(75887)

Ellen wishes to mix candy worth $1.15 per pound with candy worth $2.87 per pound to form 20 pounds of a mixture worth $1.58 per pound. How many pounds of the more expensive candy should she use?
----------
Quantity Eq::: c + e = 20 lbs.
Value Eq:: 115c +287e= 158*20 cents
========================================
Multiply thru the Quantity equation by 115:
115c + 115e = 115*20
115c + 287e = 158*20
---------------------------
Subtract and solve for "e":
172e = 43*20
Amount of expensive candey needed = 5 lbs.
=======================
cheers,
Stan H.
==========